OpenStudy (anonymous):

A culture starts with 8600 bacteria. After one hour the count is 10,000. a) Find the function that models the number of bacteria n(t) after t hours. b) Find the number of bacteria after 2 hours. c) After how many hours will the number of bacteria double?

OpenStudy (anonymous):

The form of these types of equations is \[initial~amount * 1\pm~rate~of~increase=final~amount\]

OpenStudy (anonymous):

8600*x=10,000

OpenStudy (anonymous):

What is x (our rate of increase)?

OpenStudy (anonymous):

e?

OpenStudy (anonymous):

Not this time, we're just solving for x.

OpenStudy (anonymous):

we need to put an equation together.

OpenStudy (anonymous):

We did, 8600 * rate of change = 10000

OpenStudy (anonymous):

So now you solve for the rate of change.

OpenStudy (anonymous):

Well I have n(t)= 8600e^0.1508t not using X

OpenStudy (anonymous):

That works too.

OpenStudy (anonymous):

The answer I came up with was n(t) = 8600*(1.16279)^t

OpenStudy (anonymous):

It seems you've got it.

OpenStudy (anonymous):

so plug in 2 then what about c)? about doubling it?

OpenStudy (anonymous):

Whoops, didn't see that. I'm thinking logarithms; give me a moment.

OpenStudy (anonymous):

Sorry, I don't know if I can do that one.

OpenStudy (anonymous):

b? or c?

OpenStudy (anonymous):

c. With b you just plug in 2 into your equation.

OpenStudy (anonymous):

kk! well thank you anyways

OpenStudy (anonymous):

I know how to get it with my equation; now let me try out yours.

OpenStudy (anonymous):

\[\ln 2 = 0.693\] e^0.693 * 8600 =our answer.

OpenStudy (anonymous):

It's up to you to make that exponent of e in your equation equal 0.693.

OpenStudy (ckaranja):

The number of bacteria n= initial count+bt where b is a constant n=8600+1400t